OpenStudy (anonymous):

use elimination or substitution to solve this: y=x^2 and x^2+y^2=12

OpenStudy (anonymous):

substitute x^2 in the second equation with y solve the quadratic equation.

OpenStudy (campbell_st):

use substitution substitute x^2 = y into the circle equation \[y + y^2 - 12 = 0\] which can be written as \[y^2 + y - 12 = 0\] now you have a quadratic that can be factorised... just solve for y...

OpenStudy (campbell_st):

when you have y, find x